# An “opinion reproduction number” for infodemics in a bounded-confidence content-spreading process on networks

### v1

@@ Abstract
We study the spreading dynamics of content on networks. To do this, we use a model in which content spreads through a bounded-confidence mechanism. In a bounded-confidence model (BCM) of opinion dynamics, the agents of a network have continuous-valued opinions, which they adjust when they interact with agents whose opinions are sufficiently close to theirs. The employed content-spread model introduces a twist into BCMs by using bounded confidence for the content spread itself. To study the spread of content, we define an analogue of the basic reproduction number from disease dynamics that we call an opinion reproduction number. A critical value of the opinion reproduction number indicates whether or not there is an “infodemic” (i.e., a large content-spreading cascade) of content that reflects a particular opinion. By determining this critical value, one can determine whether or not an opinion will die off or propagate widely as a cascade in a population of agents. Using configuration-model networks, we quantify the size and shape of content dissemination using a variety of summary statistics, and we illustrate how network structure and spreading-model parameters affect these statistics. We find that content spreads most widely when the agents have large expected mean degree or large receptiveness to content. When the amount of content spread only slightly exceeds the critical opinion reproduction number (i.e., the infodemic threshold), there can be longer dissemination trees than when the expected mean degree or receptiveness is larger, even though the total number of content shares is smaller.
@@
Although most content does not spread far on social media, occasionally some piece of content “goes viral” and rapidly reaches many people Goel et al. (2016); Juul and Ugander (2021). When content with misinformation or disinformation spreads widely, it has become common to expand on analogies with disease spread by stating that there is an infodemic Eysenbach (2002); Briand et al. (2021); Zielinski (2021), which is a portmanteau of the words “information” and “epidemic”. In the present paper, we examine infodemics in a model of content spread that uses a bounded-confidence mechanism. We extend the analogy between the spread of content and the spread of infectious diseases by defining an opinion reproduction number, which is inspired by the basic reproduction number of disease dynamics Kiss, Miller, and Simon (2017); Brauer, Castillo-Chavez, and Feng (2019). By examining whether or not the amount of opinion spread is below or above a critical opinion reproduction number (i.e., an infodemic threshold), one can examine whether content dies out or goes viral. This approach complements branching-process approaches that examine whether or not content spread tends to locally magnify or contract with time Gleeson et al. (2021). In our bounded-confidence model of content spread, we quantify the size and shape of content dissemination using a variety of summary statistics, and we illustrate how network structure and spreading-model parameters affect these statistics in configuration-model networks.
@@ I Introduction
Given the enormous scale and impact of human interactions on social media, it is critical to study the collective dynamics that arise in these systems Bak-Coleman et al. (2021). Content on social media can take many forms, including scientific facts, other forms of information, hyperlinks to Web pages, pictures, memes, and misinformation and disinformation Guille et al. (2013); Kümpel, Karnowski, and Keyling (2015); Muhammed T and Mathew (2022). Once created, content can then be magnified by human users, bots, and other accounts, creating an online ecosystem that is filled with much misinformation and disinformation Starbird (2019); Yang et al. (2023), even with intentions to share only accurate information Pennycook et al. (2021). Content sharing by people is ubiquitous and important for social connection, and such sharing can help create and reinforce shared understanding Baek et al. (2023). Most content does not spread to many people, but a small proportion of online content spreads very far (i.e., it can “go viral”) in large content-spreading cascades Goel et al. (2016); Juul and Ugander (2021). A variety of other social phenomena, including emotions and behaviors, can spread analogously as “social contagions” on networks Aral, Muchnik, and Sundararajan (2009); Christakis and Fowler (2013); Borge-Holthoefer et al. (2013); Yu et al. (2020); Pröllochs, Bär, and Feuerriegel (2021).
@@
When a piece of misinformation or disinformation spreads very widely, it has become reasonably common to state that there is an infodemic Eysenbach (2002); Briand et al. (2021); Zielinski (2021). The World Health Organization (WHO) gives a much more specific definition of the term “infodemic” (a portmanteau of the words “information” and “epidemic”) in the context of infectious diseases World Health Organization (2023); Calleja et al. (2021): “An infodemic is too much information including false or misleading information in digital and physical environments during a disease outbreak.” Using both the rigid WHO definition and the looser notion of particularly low-quality content going viral, the COVID-19 pandemic has also been an infodemic Zarocostas (2020); Cinelli et al. (2020); Gallotti et al. (2020). Analogously to the spread of infectious diseases “superspreader” social-media accounts have played an important role in the COVID-19 infodemic Yang et al. (2021).
@@
Developing a thorough understanding of misinformation, disinformation, and their impact requires a broad view of the problem of “fake news” that also encompasses proper understanding of misinformation and its effects Lazer et al. (2018); Watts, Rothschild, and Mobius (2021). This view needs to encompass disinformation (which is intentionally incorrect), misinformation (which is incorrect, but perhaps unintentionally), biased and misleading information (which may not be factually incorrect), and the production and amplification of such content, including by mainstream news organizations. There has been much empirical research on misinformation in areas such as computational social science and allied disciplines Muhammed T and Mathew (2022). Modeling efforts can play an important role in mitigating the harmful effects of misinformation and disinformation Rabb et al. (2022). In particular, Juul and Ugander Juul and Ugander (2021) suggested that focusing on reducing the “infectiousness” of information in concert with theoretical analysis of spreading processes may be very helpful for limiting the spread of misinformation and disinformation.
@@
One strategy to gain insight into the mechanisms that underlie observations in social-media systems is by studying mathematical models of opinion dynamics on networks Noorazar et al. (2020); Noorazar (2020); Peralta, Kertész, and Iñiguez (2022). Opinion models take a variety of forms. The nodes of a network represent agents, and the edges between agents indicate social and/or communication ties between agents. The opinions of the agents can take either discrete values (e.g., $+1$ or $-1$) or continuous values (e.g., in the interval $[-1,1]$, with $-1$ the most liberal opinion and $+1$ the most conservative opinion). When two (or more) adjacent agents interact, one or more of them updates their opinion according to some rule. One popular type of model of opinion dynamics are bounded-confidence models (BCMs) Lorenz (2007); Meng, Van Gorder, and Porter (2018); Bernardo et al. (2024), in which agents have continuous-valued opinions and interacting agents compromise their opinions by some amount if and only if their opinions are sufficiently close to each other. There have been numerous studies of BCMs, which have been generalized in many ways. Recent studies have incorporated phenomena such as media outlets with fixed opinions Brooks and Porter (2020), polyadic interactions (i.e., interactions between three or more agents) Hickok et al. (2022); Wang (2022), noise Goddard et al. (2022), asymmetric confidence intervals Bernardo, Altafini, and Vasca (2022), cost of opinion change Schawe and Hernández (2020), agents with heterogeneous activity levels Li and Porter (2023), smooth interaction kernels (in the form of sigmoidal functions) to describe how agents influence each other Brooks, Chodrow, and Porter (2023), opinion repulsion Kann and Feng (2023), homophilic adaptivity of network structure Kan, Feng, and Porter (2023), preserving connectivity of the interaction network Li, Motsch, and Weber (2020), and adaptive confidence bounds Li, Luo, and Porter (2023).
@@
Another line of modeling research is the study of spreading dynamics on networks Borge-Holthoefer et al. (2013); Porter and Gleeson (2016). Research on spreading dynamics often uses ideas from percolation theory Sahimi (2023). The study of spreading processes on networks includes research on social dynamics Lehmann and Ahn (2018), disease dynamics Pastor-Satorras et al. (2015), and how they affect each other Bedson et al. (2021). Much work on so-called “social contagions” has been inspired by the many decades of work on biological contagions, although researchers argue about whether and how much social phenomena spread in a manner that resembles disease spread Weng, Menczer, and Ahn (2013); Lerman (2016); Hui et al. (2018); Olsson and Galesic (2023); St-Onge, Hébert-Dufresne, and Allard (2024). However, despite the differences between disease spread and the spread of information and other content, it it worthwhile to explore parallels between them. Whether one is considering a disease or a piece of online content, some things spread very far (i.e., they go viral) before dissipating and others die out very quickly Juul and Ugander (2021). In some models of social phenomena, such as simplistic cascade models and anything else that takes the form of a branching process Noël et al. (2012), one can study whether the spread of content tends to magnify or contract locally as a function of time by computing branching numbers Gleeson et al. (2021), which indicate the mean number of offspring (e.g., the mean number of reposts as a function of time) of a piece of content. This is a simplification of empirical content spread on social media, where it is common to trace the spreading paths of posts, post boosting, and post commenting and quoting using dissemination trees Gomez Rodriguez, Leskovec, and Krause (2010); Guille et al. (2013); Goel et al. (2016); Kozitsin (2023). Analogously to studies of content spreading in social systems, when researchers study compartmental models of disease spread, it is traditional to calculate reproduction numbers to examine whether or not a disease will die out Brauer, Castillo-Chavez, and Feng (2019). The most standard type of reproduction number is the basic reproduction number, which measures how many infections occur, on average, from one infected node in a population in which all other nodes are susceptible to infection Kiss, Miller, and Simon (2017); Brauer, Castillo-Chavez, and Feng (2019). Compartmental models have also been employed in the study of opinion dynamics, and one can then calculate a basic reproduction number to examine whether or not content goes viral Xu et al. (2022). In the study of disease spread, branching-process theory has been used in concert with maximum-likelihood estimation to estimate basic reproduction numbers from data Blumberg and Lloyd-Smith (2013).
@@
In the present paper, we combine ideas from opinion models and percolation-inspired spreading models to study the mean “infectiousness” of content in a content-spread model that uses a bounded-confidence mechanism for the spreading. To analyze the structure of content spreading in our model, we define an opinion-dynamics analogue of the basic reproduction number from disease dynamics, and we use this opinion reproduction number and its associated critical value (which we call the “infodemic threshold”) to examine when an infodemic occurs in a network. We use generating functions to derive an analytical expression for the opinion reproduction number on configuration-model networks in the limit of infinitely many nodes. We then compare this analytical result with numerical simulations. We quantify the spread of content by calculating the total number of pieces of content, the longest adoption paths, the widths of dissemination trees, and structural virality.
@@
Our paper proceeds as follows. In section II, we give a brief introduction to bounded-confidence models of opinion dynamics. In section III, we describe a content-spread process that draws some inspiration from percolation theory and uses a bounded-confidence mechanism for spreading. In section IV, we analyze the conditions in which there is a large content-spreading cascade (i.e., an “infodemic”). To allow us to forecast whether or not an infodemic occurs in a network, we define an opinion-dynamics analogue of the basic reproduction number from disease dynamics. In section V, we quantify the total number of shared pieces of content on a finite-size network. In section VI, we compute a variety of summary statistics on dissemination trees to illustrate how network structure and spreading-model parameters affect these statistics in configuration-model networks. In section VII, we conclude and discuss our results. In section VIII, Mason writes a few words about David Campbell and wishes him a wonderful 80th birthday. The code to reproduce our numerical simulations is available at https://github.com/hzinnbrooks/bounded-confidence-spreading-process.
@@ II Bounded-confidence models on networks
We model a social network as a graph $G(V,\mathcal{E})$, where $V$ (with $|V|=N$, so $N$ is the “size” of the network) is the set of vertices (i.e., nodes) and $\mathcal{E}\subseteq V\times V$ is the set of edges. Each node represents an agent in some population, and each edge encodes a social and/or communication tie between two agents. Node $i$ is adjacent to node $j$ if there is an edge between $i$ and $j$. We assume that $G$ is simple, so it has no self-edges or multi-edges. The graph $G$ has an associated $N\times N$ adjacency matrix $A$, where $A_{ij}=1$ if $i$ is adjacent to $j$ (and vice versa) and $A_{ij}=0$ otherwise. One can consider weighted edges to model the strengths of ties between nodes. In this case, there is also a weight matrix $W$ with entries $w_{ij}\in\mathbb{R}_{\geq 0}$. If desired, one can also consider directed networks and other generalizations of graphs. See Wasserman and Faust (1994); Newman (2018); Bullo (2022); Brooks (2023) for books and reviews about networks.
Suppose that each node $i$ has a time-dependent opinion (i.e., ideology) $x_{i}(t)$, which takes continuous values on some domain that we call the opinion space. For example, the opinion space can be the real line $\mathbb{R}$, a closed and bounded interval $[a,b]\subset\mathbb{R}$, or a subset of a higher-dimensional space $\mathbb{R}^{k}$. In a BCM, the key feature is that agents are receptive only to other agents that are within a distance $c$ of their current opinion. This parameter $c$ is called the confidence bound. If $d(x_{i}(t),x_{j}(t))<c$, where $d$ is a metric on the opinion space, we say that agent $i$ is receptive to agent $j$ (and vice versa). Agents that are receptive to each other influence each other’s opinions through an update rule. It is also possible to consider BCMs with asymmetric receptiveness, but we examine only symmetric situations.

### v2

@@ Abstract
We study the spreading dynamics of content on networks. To do this, we use a model in which content spreads through a bounded-confidence mechanism. In a bounded-confidence model (BCM) of opinion dynamics, the agents of a network have continuous-valued opinions, which they adjust when they interact with agents whose opinions are sufficiently close to theirs. The employed content-spreading model introduces a twist into BCMs by using bounded confidence for the content spread itself. We define an analogue of the basic reproduction number from disease dynamics that we call an opinion reproduction number. A critical value of the opinion reproduction number indicates whether or not there is an “infodemic” (i.e., a large content-spreading cascade) of content that reflects a particular opinion. By determining this critical value, one can determine whether or not an opinion dies off or propagates widely as a cascade in a population of agents. Using configuration-model networks, we quantify the size and shape of content dissemination by calculating a variety of summary statistics, and we illustrate how network structure and spreading-model parameters affect these statistics. We find that content spreads most widely when the agents have a large expected mean degree or a large receptiveness to content. When the spreading process slightly exceeds the infodemic threshold, there can be longer dissemination trees than when the expected mean degree or receptiveness are larger, even though the total number of content shares is smaller.
@@
Although most content does not spread far on social media, occasionally some piece of content “goes viral” and rapidly reaches many people Goel et al. (2016); Juul and Ugander (2021). When content with misinformation or disinformation spreads widely, it has become common to use analogies with disease spread and state that there is an infodemic Eysenbach (2002); Briand et al. (2021); Zielinski (2021), which is a portmanteau of the words “information” and “epidemic”. In the present paper, we examine infodemics in a model of content spread that uses a bounded-confidence mechanism. We extend the analogy between the spread of content and the spread of infectious diseases by defining an opinion reproduction number, which is inspired by the basic reproduction number of disease dynamics Kiss, Miller, and Simon (2017); Brauer, Castillo-Chavez, and Feng (2019). By examining whether or not the spreading process is below or above a critical opinion reproduction number (i.e., an infodemic threshold), one can examine whether content dies out or goes viral. Our investigation complements branching-process approaches that examine whether or not content spread tends to locally magnify or contract with time Gleeson et al. (2021). In our bounded-confidence model of content spread, we quantify the size and shape of content dissemination using a variety of summary statistics, and we illustrate how network structure and spreading-model parameters affect these statistics in configuration-model networks.
@@ I Introduction
Given the enormous scale and impact of human interactions on social media, it is critical to study the collective dynamics that arise in these systems Bak-Coleman et al. (2021). Content on social media can take many forms, including scientific facts, other forms of information, hyperlinks to Web pages, pictures, memes, and misinformation and disinformation Guille et al. (2013); Kümpel, Karnowski, and Keyling (2015); Muhammed T and Mathew (2022). Once created, content can then be magnified by human users, bots, and other accounts, creating an online ecosystem that is filled with misinformation, disinformation, and “echo chambers” Friggeri et al. (2014); Starbird (2019); Yang et al. (2023); López, Pastor-Galindo, and Ruipérez-Valiente (2024); Hartmann et al. (2024), even with intentions to share only accurate information Pennycook et al. (2021). Content sharing by people is ubiquitous and important for social connection, and such sharing can help create and reinforce shared understanding Baek et al. (2023). Most content does not spread to many people, but some online content does spread very far (i.e., it can “go viral”) in large content-spreading cascades Goel et al. (2016); Juul and Ugander (2021). A variety of other social phenomena, including emotions and behaviors, can also spread as “social contagions” on networks Aral, Muchnik, and Sundararajan (2009); Christakis and Fowler (2013); Borge-Holthoefer et al. (2013); Yu et al. (2020); Pröllochs, Bär, and Feuerriegel (2021).
@@
When a piece of misinformation or disinformation spreads very widely, it has become reasonably common to state that there is an infodemic Eysenbach (2002); Briand et al. (2021); Zielinski (2021). The World Health Organization (WHO) gives a much more specific definition of the term “infodemic” (which is a portmanteau of the words “information” and “epidemic”) in the context of infectious diseases World Health Organization (2023); Calleja et al. (2021): “An infodemic is too much information including false or misleading information in digital and physical environments during a disease outbreak.” Using both the rigid WHO definition and the looser notion of particularly low-quality content going viral, it is recognized widely that the COVID-19 pandemic has had an accompanying infodemic Zarocostas (2020); Cinelli et al. (2020); Gallotti et al. (2020). Analogously to the spread of infectious diseases, “superspreader” social-media accounts have played an important role in the COVID-19 infodemic Yang et al. (2021).
@@
Developing a thorough understanding of misinformation, disinformation, and their impact requires a broad view of the problem of “fake news” that also entails proper understanding of misinformation and its effects Lazer et al. (2018); Watts, Rothschild, and Mobius (2021). This view needs to encompass disinformation (which is intentionally incorrect), misinformation (which is incorrect, but perhaps unintentionally), biased and misleading information (which may not be factually incorrect), and the production and amplification of such content, including by mainstream news organizations López, Pastor-Galindo, and Ruipérez-Valiente (2024). There has been much empirical research on misinformation in areas such as computational social science and allied disciplines Muhammed T and Mathew (2022). Importantly, modeling efforts can also play a significant role in mitigating the harmful effects of misinformation and disinformation Rabb et al. (2022). In particular, Juul and Ugander Juul and Ugander (2021) suggested that focusing on reducing the “infectiousness” of information and theoretical analyses of spreading processes may be very helpful for limiting the spread of misinformation and disinformation.
@@
One strategy to gain insight into the mechanisms that underlie observations in social-media systems is by studying mathematical models of opinion dynamics on networks Noorazar et al. (2020); Noorazar (2020); Peralta, Kertész, and Iñiguez (2022). Opinion models take a variety of forms. The nodes of a network represent agents, and the edges between agents indicate social and/or communication ties between agents. The opinions of the agents can take either discrete values (e.g., $+1$ or $-1$) or continuous values (e.g., in the interval $[-1,1]$, with $-1$ representing the most liberal opinion and $+1$ representing the most conservative opinion). When two (or more) adjacent agents interact, one or more of them updates their opinion according to some rule. One popular type of opinion model is a bounded-confidence model (BCM) Lorenz (2007); Meng, Van Gorder, and Porter (2018); Bernardo et al. (2024), in which agents have continuous-valued opinions and interacting agents compromise their opinions by some amount if and only if their opinions are sufficiently close to each other. There have been numerous studies of BCMs, which have been generalized in many ways. Recent studies have incorporated phenomena such as media outlets with fixed opinions Brooks and Porter (2020), polyadic interactions (i.e., interactions between three or more agents) Hickok et al. (2022); Wang (2022), noise Goddard et al. (2022), asymmetric confidence bounds Bernardo, Altafini, and Vasca (2022), the cost of opinion changes Schawe and Hernández (2020), agents with heterogeneous activity levels Li and Porter (2023), smooth interaction kernels (in the form of sigmoidal functions) to describe how agents influence each other Brooks, Chodrow, and Porter (2024), opinion repulsion Kann and Feng (2023), homophilic adaptivity of network structure Kan, Feng, and Porter (2023), dynamics with “no one left behind” Li, Motsch, and Weber (2020), and adaptive confidence bounds Li, Luo, and Porter (2024).
@@
Another line of modeling research is the study of spreading dynamics on networks Borge-Holthoefer et al. (2013); Porter and Gleeson (2016). Research on spreading dynamics often uses ideas from percolation theory Sahimi (2023); Malthe-Sørenssen (2024). The study of spreading processes on networks includes research on social dynamics Lehmann and Ahn (2018), disease dynamics Pastor-Satorras et al. (2015), and how they affect each other Bedson et al. (2021). Much research on so-called “social contagions” has been inspired by the rich tradition of scholarship on biological contagions, although researchers argue about whether and how much social phenomena spread in a manner that resembles disease spread Weng, Menczer, and Ahn (2013); Lerman (2016); Hui et al. (2018); Olsson and Galesic (2024); Cencetti et al. (2023); St-Onge, Hébert-Dufresne, and Allard (2024); Contreras, Cencetti, and Barrat (2024). However, despite the differences between disease spread and the spread of information and other content, it is worthwhile to explore parallels between them. Whether one is considering a disease or a piece of online content, some things spread very far before dissipating and others die out very rapidly Juul and Ugander (2021). Indeed, people even say that online content that spreads very far has “gone viral”.
In some models of social phenomena, such as simplistic cascade models and any other models in the form of a branching process Noël et al. (2012), one can study whether the spread of content tends to magnify or contract locally as a function of time by computing branching numbers Gleeson et al. (2021), which indicate the mean number of offspring (e.g., the mean number of reposts as a function of time) of a piece of content. This is a simplification of empirical content spread on social media, where it is common to trace the spreading paths of posts, post boosting, and post commenting and quoting using dissemination trees Gomez Rodriguez, Leskovec, and Krause (2010); Guille et al. (2013); Goel et al. (2016); Kozitsin (2023). There are also other ways to calculate local magnification and contraction in spreading processes Baños, Borge-Holthoefer, and Moreno (2013). Analogously to studies of content spread in social systems, when researchers study compartmental models of disease spread, it is traditional to calculate reproduction numbers to examine whether or not a disease dies out Brauer, Castillo-Chavez, and Feng (2019). The most standard type of reproduction number is the basic reproduction number, which measures how many infections occur, on average, from one infected node in a population in which all other nodes are susceptible to infection Kiss, Miller, and Simon (2017); Brauer, Castillo-Chavez, and Feng (2019). Compartmental models have also been employed in studies of opinion dynamics, and one can then calculate a basic reproduction number to examine whether or not content goes viral Xu et al. (2022). In the study of disease spread, branching-process theory has also been used in concert with maximum-likelihood estimation to estimate basic reproduction numbers from data Blumberg and Lloyd-Smith (2013).
In the present paper, we combine ideas from opinion models and percolation-inspired spreading models to study the mean “infectiousness” of content in a model that uses a bounded-confidence mechanism for content spread [^1] A complementary approach is to modify opinion models by incorporating content Wang, Chen, and Zhao (2024).. In our content-spreading model, we use a standard bounded-confidence mechanism to determine whether or not an agent is receptive to content. If an agent is receptive, it spreads content without modifying it, rather than compromising its opinion as in a standard BCM. Our model takes the form of a threshold model Lehmann and Ahn (2018) and has an opinion-update rule that is reminiscent of bond percolation Sahimi (2023); Dorogovtsev, Goltsev, and Mendes (2008); Newman (2018). (It is also related to the independent-cascade model and its generalizations Kempe, Kleinberg, and Tardos (2003).) In standard bond percolation, an edge of a network is preserved with probability $p$ and is removed with probability $1-p$. In our content-spreading model, the probability of “preserving” an edge between agents $i$ and $j$ corresponds to the spread of content between those two agents. However, in contrast to classical bond-percolation models, this probability is not fixed in our model. Instead, it depends on the opinions of agents $i$ and $j$ and and on whether or not those two agents are adjacent to other agents that previously shared the content.
@@
To analyze the structure of content spread in our model, we define an opinion-dynamics analogue of the basic reproduction number from disease dynamics, and we use this opinion reproduction number and its associated critical value (which we call the infodemic threshold) to examine when an infodemic occurs in a network. We use generating functions to derive an analytical expression for the opinion reproduction number on configuration-model networks in the limit of infinitely many nodes. We then compare this analytical result with numerical simulations. We quantify the spread of content by calculating a variety of summary statistics — the total number of shared pieces of content, the longest adoption paths, the widths of dissemination trees, and structural virality — that were employed previously by other researchers Vosoughi, Roy, and Aral (2018); Juul and Ugander (2021).
@@
Our paper proceeds as follows. In Section II, we give a brief introduction to bounded-confidence models of opinion dynamics. In Section III, we describe a content-spreading process that draws inspiration from percolation theory and uses a bounded-confidence mechanism for spreading. In Section IV, we analyze the conditions that determine when we expect a large content-spreading cascade (i.e., an “infodemic”). To allow us to forecast whether or not an infodemic occurs in a network, we define an opinion-dynamics analogue of the basic reproduction number from disease dynamics. In Section V, we quantify the total number of shared pieces of content on a finite-size network. In Section VI, we compute a variety of summary statistics to describe dissemination trees, and we illustrate how network structure and spreading-model parameters affect these statistics in configuration-model networks. In Section VII, we conclude and discuss our results. In Appendix A, Mason writes a few words about David Campbell and wishes him a wonderful 80th birthday. The code to reproduce our numerical simulations is available at https://github.com/hzinnbrooks/bounded-confidence-spreading-process.
@@ II Bounded-confidence models on networks
We model a social network as a graph $G(V,\mathcal{E})$, where $V$ (with $|V|=N$, so $N$ is the “size” of the network) is the set of vertices (i.e., nodes) and $\mathcal{E}\subseteq V\times V$ is the set of edges. Each node represents an agent in some population, and each edge encodes a social and/or communication tie between two agents. We suppose that $G$ is undirected and unweighted. Node $i$ is adjacent to node $j$ (and vice versa) if there is an edge between $i$ and $j$. We assume that $G$ is simple, so it has no self-edges or multi-edges. The graph $G$ has an associated $N\times N$ adjacency matrix $A$, where $A_{ij}=1$ if $i$ is adjacent to $j$ and $A_{ij}=0$ otherwise. If desired, one can consider directed networks, weighted networks, and other generalizations of graphs. See Refs. [Wasserman and Faust, 1994; Newman, 2018; Bullo, 2022; Brooks, 2023] for books and reviews about networks.
Suppose that each node $i$ has a time-dependent opinion $x_{i}(t)$, which takes continuous values on a domain that we call the opinion space. For example, the opinion space can be the real line $\mathbb{R}$, a closed and bounded interval $[a,b]\subset\mathbb{R}$, or a subset of a higher-dimensional space $\mathbb{R}^{k}$. In a BCM, agents are receptive only to other agents that are within a distance $c$ of their current opinion. More precisely, we say that agent $i$ is receptive to agent $j$ (and vice versa) if $d(x_{i}(t),x_{j}(t))<c$, where $d$ is a metric on the opinion space. The parameter $c$ is called the confidence bound. Agents that are receptive to each other influence each other’s opinions through an update rule. It is also possible to consider BCMs with asymmetric receptiveness, but we examine only symmetric situations.